OpenStudy (anonymous):

2/3 = 4j/6 has to show work..

OpenStudy (nirmalnema):

do u want to find the value of J

OpenStudy (nirmalnema):

2/3 = 4j/6 multiply both side by 6 you will get 2*2 / = 4j 4/4= j j= 1

OpenStudy (nirmalnema):

@ayeissel did u understand